OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14

jhonyy9 (jhonyy9):

how you need solve it by elimination or substitution method ?

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

you need to solve this system for get the x value ,yes

jhonyy9 (jhonyy9):

ok try multiplie first equation by 5 and the second by -3 what will get ?

jhonyy9 (jhonyy9):

4x + 3y = 10 *5 4x + 5y = 14 *(-3) 20x +15y = 50 -12x -15y = - 42 ------------------- add term by term 8x 0 = 8 8x = 8 x = 8/8 x = 1

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yes thank you

jhonyy9 (jhonyy9):

yw good luck

OpenStudy (anonymous):

x=1